OpenStudy (anonymous):

Can someone check my answer? Thanks! :) Select the graph and end behavior description of f(x) = 3 + x6. Graph is a flattened parabola with vertex at 0, 3 and opens down. This is an even function. The end behavior of the graph will decrease to the right and to the left.

jimthompson5910 (jim_thompson5910):

the leading term is x^6 it is positive, so this means that the graph must open upwards

OpenStudy (anonymous):

Oh!! Would the vertex be at 0, 3 or 0, -3?

jimthompson5910 (jim_thompson5910):

well when x = 0, what is y?

jimthompson5910 (jim_thompson5910):

to find out, plug it into the function and evaluate

OpenStudy (anonymous):

I'm not 100% sure how to do that. :( Could you teach me?

jimthompson5910 (jim_thompson5910):

the function is f(x) = 3 + x^6

jimthompson5910 (jim_thompson5910):

plug in x = 0 to get f(x) = 3 + x^6 f(0) = 3 + (0)^6 f(0) = ???

OpenStudy (anonymous):

positive 3 so the vertex would be 0, 3

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

(0,3) actually, but you got the idea

OpenStudy (anonymous):

thank you thank you! :) Oh! (0,3)

jimthompson5910 (jim_thompson5910):

you're welcome and the end behavior would be seen if you graphed it or tried to visualize it in your mind